Let G be a subgroup of GL2 (Z4) defined by the set {[m b ,0 1}] such that b ∈ Z4 and m=±1. Show that G is isomorphic to a known group of order 8?


1
Expert's answer
2021-05-26T15:43:01-0400

Given

G be a subgroup of GL(Z4

Define as



"G={\begin{bmatrix}
 m & b \\
 0 & 1
\end{bmatrix}}"



Such that b belong to Zand m=±1.

Now we Show that G is isomorphic to a known group of order 8.

Let known group of order 8 is Z× Z4


Show to isomoriphic ,then we will be show one one , onto , and morphism.


Define a map

f: G to Z× Z4



"f(\begin{bmatrix}
 m & b \\
 0 & 1
\end{bmatrix})"

=(m,b)



"One one"

If we take any 2 by 2 matrix such as above then

If both element of 1 st row is same then mapping is same.

It show f is one one.

"Onto"

Each elements belong to Z×Z4


There exist a matrix in G .then we say that it is onto.


"Homomophism"

f(aA+bB)=af(A)+bf(B)

Because we know that matrix alsway show the prppery of homomorphism.


So finally we say that G os isomorphic to known group of order 8.